OpenStudy (anonymous):

Help in math 2B for connections and need to pass!!

OpenStudy (anonymous):

\(\Large\rm\color{orange}{Welcome~to~OpenStudy!}\) What do you need help with?

OpenStudy (anonymous):

\[2x +4=6x -20\]

OpenStudy (anonymous):

need to find value of x

OpenStudy (anonymous):

24 = 8x 3 = x

OpenStudy (anonymous):

you have to add like terms and then you use pemdas

OpenStudy (owlcoffee):

That is what we call a "first degree equation", why?. Because the highest exponent of the variable is "1", you'll learn that further in algebra, let's just keep it at that. Whenever you solve an equation, you "manipulate it" by adding, sustracting, dividing or multiplying any quatity, variable or even expressions some times. So, let's take that equation of yours: \[2x+4=6x-20\] The main idea is to "manipulate" the equation and leave the x alone on either sides of the "=" sign. So, let's do that: We will begin by sustracting "4" on either sides of the equation: \[2x+4-4=6x-20-4\] and operating: \[2x=6x-24\] Now, let's sustract "6x" on both sides as well: \[2x-6x=6x-6x-24\] and operate: \[-4x=-24\] And to isolate that "x", we will divide both sides of the equation by "-4": \[\frac{ -4x }{ -4 }=\frac{ -24 }{ -4 }\] and operating: \[x=\frac{ -24 }{ -4 }\] and "-24/-4" is equal to 6, so that means: \[x=6\] And there you go.

OpenStudy (anonymous):

wait yeah sorry I added 2x instead of subtracting so itd be 24 = 4x 6 = x my bad